# LLM as GNN: Graph Vocabulary Learning for Text-Attributed Graph Foundation Models

### v1

@@ 6 Conclusion
7 Limitations
The proposed PromptGFM framework introduces a promising paradigm for building a Graph Foundation Model (GFM) in Text-Attributed Graphs (TAGs). However, like any novel framework, it has limitations that should be acknowledged for future improvements. First, while we establish a foundation model interface, incorporating a large number of datasets may negatively impact transferability, particularly in cross-domain scenarios. A more stable knowledge transfer mechanism is needed to mitigate domain shifts and preserving generalization. Second, our prompt-based GNN processes each graph independently, disregarding dataset imbalances and cross-graph semantic similarities. One solution is to explore cross-graph semantic alignment and attentive learning for better knowledge integration across heterogeneous graph datasets. Third, our approach remains constrained to text-attributed graphs, leaving a gap in developing a fully text-free graph foundation model. These limitations highlight potential directions for future research in building more efficient and comprehensive GFMs.
@@ 7 Limitations
  • §A provides Data Descriptions used in our experiments.
  • §B illustrates more details about Baselines employed for comparison.
  • §C shows systematic comparison between Embedding-based GNNs and Prompt-based GNNs.
  • §D analyzes Prompt Design in our PromptGFM framework.
  • §E shows more Implementation Details.
  • §F reports the Hyperparameter Sensitivity for prompt-based GNNs.
  • §G presents the Case Study to offer deeper insights of our research.
@@ A.1 Computer Science Datasets
$\bullet$ Cora (McCallum et al., 2000). The dataset includes a citation network consisting of 2,708 scientific publications in the field of machine learning, categorized into 7 classes based on their research topics. Nodes represent individual papers, and edges denote citation links between them, totaling 10,858 connections. Each paper is associated with its title and abstract.
$\bullet$ Citeseer (Giles et al., 1998). This dataset introduces a citation network comprising 3,327 scientific publications, categorized into 6 classes, including Agents, Artificial Intelligence, Database, Information Retrieval, Machine Learning, and Human-Computer Interaction. Nodes correspond to documents, and edges represent citation relationships between them, amounting to 9,464 links.
$\bullet$ obgn-arxiv (Hu et al., 2020). The ogbn-arxiv dataset is part of the Open Graph Benchmark and consists of a directed citation graph of 169,343 arXiv papers, categorized into 40 subject areas. Nodes represent individual papers, and edges indicate citation relationships, totaling over 2.3 million connections. Each paper includes textual data from its title and abstract.
$\bullet$ WikiCS (Mernyei and Cangea, 2020). The WikiCS dataset is a benchmark dataset derived from Wikipedia, designed for evaluating GNNs. It comprises a citation network with 11,701 nodes representing computer science articles and 431,726 edges corresponding to hyperlinks between them. The dataset features 10 classes corresponding to branches of computer science, with very high connectivity.
@@ A.2 E-commerce Datasets
$\bullet$ History (Yan et al., 2023). The History dataset, extracted from the Amazon-Books dataset, includes items categorized under the second-level label "History". It comprises 41,511 nodes and 503,180 edges, where each node represents a book, and edges indicate frequent co-purchases or co-views between books. This dataset incorporates the title and description of each book as text attributes of the node. The classification task involves assigning books to 12 distinct categories.
$\bullet$ Photo (Yan et al., 2023). The Photo dataset, derived from the Amazon-Electronics dataset, consists of 48,362 nodes and 873,782 edges, forming a dense network that reflects user purchasing behavior. Each node represents an electronic product, while edges indicate frequent co-purchases or co-views between items. For textual attributes, user reviews are incorporated, prioritizing the most upvoted review when available; otherwise, a randomly selected review is used. The classification task involves categorizing electronic products into 12 distinct categories.
@@ A.3 Biomedical Datasets
$\bullet$ PubMed (Sen et al., 2008). PubMed is a citation network of 19,717 scientific publications from the PubMed database pertaining to diabetes, classified into 3 classes: experimental induced diabetes, type 1 diabetes, and type 2 diabetes. Nodes are research papers, and edges signify citation links, amounting to 88,648 connections. This dataset is used for large-scale graph representation learning and evaluating algorithms in the biomedical domain.
@@ Appendix B Baselines
We provide detailed information on the baseline models, categorized into: (1) Graph-agnostic methods, (2) GNN-based methods, (3) BERT-based methods, and (4) GNN-LLM integration methods.
@@
$\bullet$ GCN  (Kipf and Welling, 2017). This model introduces a neural network architecture that generalizes convolution operations to graph-structured data, enabling effective semi-supervised learning by aggregating feature information from a node’s local neighborhood.
$\bullet$ GAT  (Velickovic et al., 2017). This method incorporates attention mechanisms into graph neural networks, allowing nodes to assign different importance weights to their neighbors during feature aggregation, which enhances performance by focusing on the most relevant connections.
$\bullet$ SAGE  (Hamilton et al., 2017). GraphSAGE is an inductive representation learning framework on large graphs; it generates node embeddings by sampling and aggregating features from a node’s local neighborhood, facilitating generalization to unseen nodes or graphs.
$\bullet$ ReVGNN  (Li et al., 2021). This method includes a recurrent graph neural network tailored for dynamic graphs, capturing temporal dependencies by updating node representations as events occur over time, which is crucial for modeling evolving graph structures.
$\bullet$ SGFormer  (Wu et al., 2023). This work introduces a transformer-based architecture designed for graph data, integrating spectral graph theory into the transformer framework. It aims to capture both local and global graph structures efficiently by incorporating spectral filters, enhancing the model’s ability to learn complex graph representations.
$\bullet$ NodeFormer  (Wu et al., 2022). This framework presents a scalable graph transformer model that utilizes a randomized attention mechanism to approximate full attention on graphs. By reducing computational complexity, it enables efficient learning on large-scale graphs while preserving the expressiveness of transformer architectures.
B.3 BERT-based methods.
$\bullet$ BERT  (Devlin et al., 2019). This work introduces a deep learning model that understands language context by processing text bidirectionally. It is pre-trained using masked language modeling (MLM) and next sentence prediction (NSP) to learn rich linguistic features. BERT can be fine-tuned for various NLP tasks, achieving state-of-the-art performance.
$\bullet$ Sentence-BERT  (Reimers and Gurevych, 2019). This model is a modification of BERT designed for sentence embeddings, allowing efficient comparison of semantic similarity. It uses a siamese or triplet network structure to generate fixed-size vector representations, making tasks like sentence similarity, clustering, and retrieval significantly faster. Unlike BERT, which requires computationally expensive cross-encoding, Sentence-BERT enables quick and effective comparisons using cosine similarity.
@@
$\bullet$ LLaGA  (Chen et al., 2024a). This model effectively integrates LLM capabilities to handle the complexities of graph-structured data. It transforms graph nodes into structure-aware sequences and maps them into token embedding space using a specialized projector. LLaGA excels in generalization and interpretability, performing strongly across various datasets and tasks. It also supports zero-shot learning, making it highly adaptable for unseen datasets.
$\bullet$ OFA  (Liu et al., 2024a). This paper proposes a framework that handles various graph classification tasks across different domains using a single model. It introduces the nodes-of-interest (NOI) subgraph mechanism to standardize different tasks with a single task representation. Additionally, a novel graph prompting paradigm to leverage in-context learning and apply the same architecture across diverse graph classification tasks, achieving generalization across multiple domains.
$\bullet$ GraphPrompter  (Liu et al., 2024b). This work introduces a novel framework designed to align graph with LLMs via soft prompts. Specifically, it adopts GNNs to capture graph structure and leverages an LLM to interpret the textual information at the node level. By prompt tuning, this approach demonstrates the potential of LLMs to effectively interpret graph structures, combining both semantic and structural insights for improved graph learning tasks.
$\bullet$ ENGINE  (Zhu et al., 2024). This paper proposes a parameter- and memory-efficient fine-tuning method for textual graphs by using LLMs as encoders. It combines the LLMs and GNNs through a tunable GNN-based side structure, called G-Ladder, alongside each LLM layer, effectively reducing training costs without compromising performance.
Figure 6: Graph understanding module via prompt-based GNNs. We prompt LLMs to achieve fine-grained reproduction of traditional GNN workflow, refining verbose textual representations into concise yet meaningful ones. In the prompt, neighbor sampling (see Equation 3) is highlighted in purple, the aggregation-update mechanism (see Equation 4) in blue, and the optimization in red.
@@
Table 6 illustrates the details of the graph understanding module, with the prompt-based GNN as its core component. To ensure clarity, we systematically compare traditional embedding-based GNNs with our proposed prompt-based GNN, emphasizing its advantages.
@@ Appendix E Implementation Details
We provide further information for reproduction. In the graph understanding module, we selected the number of layers for the prompt-based GNN from $\{1,2,3,4\}$. We randomly sampled 30% of the first-order neighbors during neighborhood sampling, capping the maximum number of sampled nodes at 20 to reduce computational cost and prevent overfitting. In the graph inference module, we fine-tuned the LLM with a learning rate of 3e-4 and a batch size of 4. To mitigate potential biases introduced by task-specific prompts, we designed a prompt pool for each task requirement and randomly selected prompts during instruction construction to enhance robustness. We employ 10-fold cross-validation and report average results across all folds. We employed a standard early-stopping strategy during training: if the performance metric on the validation set did not improve over a fixed number of consecutive epochs (determined based on the dataset), we halted training to prevent overfitting. For other hyperparameters of the compared methods, we referred to the original papers and carefully tuned them to suit each dataset. In terms of the TAG setting, we utilize the textual features for initialization in all embedding-based models.
Appendix F Hyperparameter Sensitivity
Figure 7: Impact of varying prompt-based GNN layers on node classification performance.
We explore the impact of the number of layers in our prompt-based GNN. As shown in Figure 7, we can observe that PromptGFM progressively improves as the layer increases due to its ability to capture broader context and higher-order relationships over the graph. However, after a certain point, further stacking layers results in diminishing returns or even performance degradation due to over-smoothing, where node representations become indistinguishable within their local structures. This trend is consistent with traditional GNNs. Optimal performance is achieved with 3-layer GNN for Cora, while Citeseer reaches its best results with 2 layers. This analysis suggests that textual representations can be propagated over the graph similarly to numerical embeddings, effectively capturing semantic and structural information simultaneously.
Appendix G Case Study
G.1 Textual Representations in Prompt-based GNNs
@@ G.1.2 Case B
G.2 Language-based IDs vs. Key Words
@@
Table 9: The comparison between language-based IDs generated by prompt-based GNNs as part of the graph vocabulary and keywords extracted from the original papers in citation datasets. To prevent data leakage, we exclude these keywords from model training. The observed similarity confirms that our prompt-based GNN effectively captures the essential semantics of nodes. Furthermore, our language-based IDs encode richer information than keywords by preserving high-order structural signals.

### v2

@@ 6 Conclusion
$\bullet$ §A provides Data Descriptions used in our experiments.
$\bullet$ §B illustrates more details about Baselines employed for comparison.
$\bullet$ §C shows comparison between Embedding-based and Prompt-based GNNs.
$\bullet$ §D analyzes Prompt Design in our PromptGFM framework.
$\bullet$ §E shows more Implementation Details.
$\bullet$ §F provides a Prompt Length Limitation Discussion.
$\bullet$ §G investigates Permutation Sensitivity in prompt-based GNNs.
$\bullet$ §H presents the Case Study to offer deeper insights of our research.
@@ A.1 Computer Science Datasets
$\bullet$ Cora [32]. The dataset includes a citation network consisting of 2,708 scientific publications in the field of machine learning, categorized into 7 classes based on their research topics. Nodes represent individual papers, and edges denote citation links between them, totaling 10,858 connections. Each paper is associated with its title and abstract.
$\bullet$ Citeseer [13]. This dataset introduces a citation network comprising 3,327 scientific publications, categorized into 6 classes, including Agents, Artificial Intelligence, Database, Information Retrieval, Machine Learning, and Human-Computer Interaction. Nodes correspond to documents, and edges represent citation relationships between them, amounting to 9,464 links.
$\bullet$ obgn-arxiv [18]. The ogbn-arxiv dataset is part of the Open Graph Benchmark and consists of a directed citation graph of 169,343 arXiv papers, categorized into 40 subject areas. Nodes represent individual papers, and edges indicate citation relationships, totaling over 2.3 million connections. Each paper includes textual data from its title and abstract.
$\bullet$ WikiCS [33]. The WikiCS dataset is a benchmark dataset derived from Wikipedia, designed for evaluating GNNs. It comprises a citation network with 11,701 nodes representing computer science articles and 431,726 edges corresponding to hyperlinks between them. The dataset features 10 classes corresponding to branches of computer science, with very high connectivity.
@@ A.2 E-commerce Datasets
$\bullet$ History [57]. The History dataset, extracted from the Amazon-Books dataset, includes items categorized under the second-level label "History". It comprises 41,511 nodes and 503,180 edges, where each node represents a book, and edges indicate frequent co-purchases or co-views between books. This dataset incorporates the title and description of each book as text attributes of the node. The classification task involves assigning books to 12 distinct categories.
$\bullet$ Photo [57]. The Photo dataset, derived from the Amazon-Electronics dataset, consists of 48,362 nodes and 873,782 edges, forming a dense network that reflects user purchasing behavior. Each node represents an electronic product, while edges indicate frequent co-purchases or co-views between items. For textual attributes, user reviews are incorporated, prioritizing the most upvoted review when available; otherwise, a randomly selected review is used. The classification task involves categorizing electronic products into 12 distinct categories.
@@ A.3 Biomedical Datasets
$\bullet$ PubMed [42]. PubMed is a citation network of 19,717 scientific publications from the PubMed database pertaining to diabetes, classified into 3 classes: experimental induced diabetes, type 1 diabetes, and type 2 diabetes. Nodes are research papers, and edges signify citation links, amounting to 88,648 connections. This dataset is used for large-scale graph representation learning and evaluating algorithms in the biomedical domain.
@@ Appendix B Baselines
We provide detailed information on the baseline models, categorized into: (1) Graph-agnostic methods, (2) GNN-based methods, (3) LM-only methods, and (4) GNN-LLM integration methods.
@@
$\bullet$ GCN  [23]. This model introduces a neural network architecture that generalizes convolution operations to graph-structured data, enabling effective semi-supervised learning by aggregating feature information from a node’s local neighborhood.
$\bullet$ GAT  [46]. This method incorporates attention mechanisms into graph neural networks, allowing nodes to assign different importance weights to their neighbors during feature aggregation, which enhances performance by focusing on the most relevant connections.
$\bullet$ SAGE  [16]. GraphSAGE is an inductive representation learning framework on large graphs; it generates node embeddings by sampling and aggregating features from a node’s local neighborhood, facilitating generalization to unseen nodes or graphs.
$\bullet$ ReVGNN  [24]. This method includes a recurrent graph neural network tailored for dynamic graphs, capturing temporal dependencies by updating node representations as events occur over time, which is crucial for modeling evolving graph structures.
$\bullet$ SGFormer  [54]. This work introduces a transformer-based architecture designed for graph data, integrating spectral graph theory into the transformer framework. It aims to capture both local and global graph structures efficiently by incorporating spectral filters, enhancing the model’s ability to learn complex graph representations.
$\bullet$ NodeFormer  [53]. This framework presents a scalable graph transformer model that utilizes a randomized attention mechanism to approximate full attention on graphs. By reducing computational complexity, it enables efficient learning on large-scale graphs while preserving the expressiveness of transformer architectures.
B.3 LM-only methods.
$\bullet$ BERT  [10]. This work introduces a deep learning model that understands language context by processing text bidirectionally. It is pre-trained using masked language modeling (MLM) and next sentence prediction (NSP) to learn rich linguistic features. BERT can be fine-tuned for various NLP tasks, achieving state-of-the-art performance.
$\bullet$ Sentence-BERT  [40]. This model is a modification of BERT designed for sentence embeddings, allowing efficient comparison of semantic similarity. It uses a siamese or triplet network structure to generate fixed-size vector representations, making tasks like sentence similarity, clustering, and retrieval significantly faster. Unlike BERT, which requires computationally expensive cross-encoding, Sentence-BERT enables quick and effective comparisons using cosine similarity.
$\bullet$ LLaMA3  [27]. LLaMA3-8B is an open-weight transformer model with 8 billion parameters, part of Meta’s third-generation LLaMA series. It is trained on 15T tokens from a diverse mix of web, code, and academic data, using a decoder-only architecture with a context window of 8,192 tokens. LLaMA3-8B achieves strong performance on a wide range of benchmarks, particularly in reasoning, multilingual tasks, and instruction following. Compared to earlier versions, it benefits from improved training stability, data quality, and scaling efficiency, making it a competitive foundation model for downstream adaptation and alignment.
@@
$\bullet$ LLaGA  [6]. This model effectively integrates LLM capabilities to handle the complexities of graph-structured data. It transforms graph nodes into structure-aware sequences and maps them into token embedding space using a specialized projector. LLaGA excels in generalization and interpretability, performing strongly across various datasets and tasks. It also supports zero-shot learning, making it highly adaptable for unseen datasets.
$\bullet$ OFA  [27]. This paper proposes a framework that handles various graph classification tasks across different domains using a single model. It introduces the nodes-of-interest (NOI) subgraph mechanism to standardize different tasks with a single task representation. Additionally, a novel graph prompting paradigm to leverage in-context learning and apply the same architecture across diverse graph classification tasks, achieving generalization across multiple domains.
$\bullet$ GraphPrompter  [29]. This work introduces a novel framework designed to align graph with LLMs via soft prompts. Specifically, it adopts GNNs to capture graph structure and leverages an LLM to interpret the textual information at the node level. By prompt tuning, this approach demonstrates the potential of LLMs to effectively interpret graph structures, combining both semantic and structural insights for improved graph learning tasks.
$\bullet$ ENGINE  [62]. This paper proposes a parameter- and memory-efficient fine-tuning method for textual graphs by using LLMs as encoders. It combines the LLMs and GNNs through a tunable GNN-based side structure, called G-Ladder, alongside each LLM layer, effectively reducing training costs without compromising performance.
Figure 7: Graph understanding module via prompt-based GNNs. We prompt LLMs to achieve fine-grained reproduction of traditional GNN workflow, refining verbose textual representations into concise yet meaningful ones. In the prompt, neighbor sampling (see Equation 3) is highlighted in darkred, the aggregation-update mechanism (see Equation 4) in blue, and the optimization in red.
@@
Figure 7 illustrates the details of the graph understanding module, with the prompt-based GNN as its core component. To ensure clarity, we systematically compare traditional embedding-based GNNs with our proposed prompt-based GNN, emphasizing its advantages.
@@ Appendix E Implementation Details
We provide additional details to facilitate reproducibility. In the graph understanding module, the number of layers in the prompt-based GNN is selected from ${1,2,3,4}$, and the first-order neighbor sampling ratio from ${0.3,0.6,1.0}$. In the graph inference module, we fine-tune the LLM with a learning rate of 3e-4 and a batch size of 4. To mitigate potential biases introduced by task-specific prompts, we designed a prompt pool for each task requirement and randomly selected prompts during instruction construction to enhance robustness. We employed a standard early-stopping strategy during training: if the performance metric on the validation set did not improve over a fixed number of consecutive epochs (determined based on the dataset), we halted training to prevent overfitting. For other hyperparameters of the compared methods, we referred to the original papers and carefully tuned them to suit each dataset.
Appendix F Prompt Length Limitation Discussion
In dense TAGs, high-degree nodes pose a challenge for prompt-based GNNs due to potentially excessive prompt lengths, which can exceed the input limits of current LLMs and impair model efficiency. To address this issue, we analyze in the following aspects that balances efficiency, scalability, and representation quality.
F.1 Neighbor Sampling in Scalable GNNs
Following standard GNN practices, we sample a fixed number of neighbors per node to reduce prompt length and improve robustness. As shown in Table 11, varying the sampling ratio across three datasets reveals that selecting 60% of neighbors yields the best performance, confirming its effectiveness. Meanwhile, limiting the maximum number of sampled neighbors is a common practice in scalable GNNs. For example, GraphSAGE samples 25 and 10 neighbors at the first and second hops, respectively, while PinSAGE selects from $\{10,20,50\}$. Inspired by these practices, we choose from $\{10,20,30,40\}$ and empirically set 20 as the cap, ensuring prompts remain within LLM input limits while preserving sufficient structural context.
F.2 Token Capacity of LLMs
Our implementation leverages GPT-4o mini, which supports up to 128k input tokens per call. This capacity allows a single prompt to encode information from over 300 neighbors, given that each neighbor’s textual summaries occupies approximately 400 tokens on average (for the first round of message propagation). As shown in Table 12, only a small fraction of nodes (<0.25%) in a few datasets exceed this limit. For these rare cases, we have applied neighbor sampling and truncation in Section F.1, effectively bounding prompt length without degrading representation quality.
F.3 Prompt Design
In our prompt-based GNN framework, we design prompts that guide the LLM to aggregate information from a node’s sampled neighbors and produce a concise yet expressive representation. This process is inherently iterative: with each layer, the prompt becomes more compressed as semantic and structural information is distilled into fewer tokens. Empirically, the average prompt length decreases from approximately 400 tokens in the first layer, to  125 tokens in the second, and eventually to  10 tokens per node on the Cora dataset.
Importantly, each node is processed independently, and its prompt length depends solely on its local neighborhood, i.e., its degree, rather than on global graph size. This localized design ensures scalability, as prompt lengths remain bounded even in large graphs, allowing our approach to maintain efficiency without sacrificing representational capacity.
Table 11: The performance with varying neighbor sampling ratios acorss three datasets.
Sampling Ratio 100% 60% 30%
Cora 81.36 89.35 88.36
Citeseer 79.34 82.23 81.13
PubMed 90.23 91.12 89.34
Table 12: Analysis of High-Degree Nodes and Average Connectivity
Metric Cora Citeseer PubMed Ogbn-arxiv History Photo WikiCS
Avg. Degree 4.01 2.84 13.77 36.90 18.07 12.11 4.50
High-degree (>300) Nodes 0 0 0 0.20% 0.25% 0.19% 0
Appendix G Permutation Sensitivity
We evaluate the sensitivity of our prompt-based GNNs (using Flan-T5) to input order by evaluating two variants: (a) shuffle nodes, which randomizes the order of neighbors, and (b) shuffle tokens, which injects position-specific tokens to nodes and shuffle them instead. As illustrated in Table 13, regardless of the shuffle method, both variants result in only minor fluctuations in node classification accuracy, indicating that prompt-based GNNs with the subsequent fine-tuning are robust to input permutation and maintain stable performance.
Table 13: Permutation sensitivity of prompt-based GNNs.
      Variant       Cora       Citeseer
      shuffle nodes       90.64 $\pm$1.12       83.79 $\pm$ 0.89
      shuffle tokens       90.55 $\pm$1.03       84.28 $\pm$1.06
Appendix H Case Study
H.1 Textual Representations in Prompt-based GNNs
@@ H.1.2 Case B
H.2 Language-based IDs vs. Key Words
@@
Table 14: The comparison between language-based IDs generated by prompt-based GNNs as part of the graph vocabulary and keywords extracted from the original papers in citation datasets. To prevent data leakage, we exclude these keywords from model training. The observed similarity confirms that our prompt-based GNN effectively captures the essential semantics of nodes. Furthermore, our language-based IDs encode richer information than keywords by preserving high-order signals.
@@
H.3 Impact Statement
We highlight the following concrete cases to better demonstrate the broad applicability of our approach.
$\bullet$ In recommender systems, different scenarios involve distinct graph structures and scenario-specific semantics. Instead of training scenario-specific models, our foundation model could act as a general-purpose recommendation backbone that unifies multiple scenarios in text space via a co-training paradigm. This enables LLMs to capture rich and transferable semantic relationships across scenarios (e.g., user preferences, item affinities, contextual behaviors). Specifically, (a) our foundation model alleviates cross-dataset inconsistencies and facilitates rapid adaptation to new scenarios and tasks. (b) it is particularly effective in cold-start issues and long-tail recommendation challenges, where existing methods often struggle due to rigid architectures and heavy dependence on labeled data.
$\bullet$ In biomedical knowledge graphs, entities (e.g., genes, proteins, drugs, diseases) and their complex relations (e.g., inhibition, expression, association) form semantically rich graphs. Our approach achieves cross-resource knowledge integration (e.g., DrugBank, CTD, DisGeNET) and supports inductive reasoning across biomedical subdomains, enabling fast adaptation to new tasks and unseen relation types with minimal supervision. This is particularly valuable in biomedical settings, where knowledge resources are often fragmented, relation annotations are sparse, and entity semantics vary significantly across datasets. In this context, our approach is well suited for low-resource scenarios such as rare disease research, novel drug discovery, and cross-database knowledge transfer.